OpenStudy (unofficialllyy):

I really need help with this question. A Web music store offers two versions of a popular song. The size of the standard version is 2.5 megabytes (MB). The size of the high-quality version is 4.1 MB. Yesterday, the high-quality version was downloaded four times as often as the standard version. The total size downloaded for the two versions was 4536 MB. How many downloads of the standard version were there?

OpenStudy (unofficialllyy):

I just need help setting up the equation

OpenStudy (unofficialllyy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Let x = number of standard version songs downloaded y = number of high-quality version songs downloaded

jimthompson5910 (jim_thompson5910):

`Yesterday, the high-quality version was downloaded four times as often as the standard version` so we know that (number of high quality) = 4*(number of standard) y = 4x

jimthompson5910 (jim_thompson5910):

`The size of the standard version is 2.5 megabytes (MB). The size of the high-quality version is 4.1 MB. ... The total size downloaded for the two versions was 4536 MB.` so the other equation is 2.5x+4.1y = 4536

OpenStudy (ghostgate):

Alright, in setting up the equation we would do this: "A Web music store offers two versions of a popular song. The size of the standard version is 2.5 megabytes (MB). The size of the high-quality version is 4.1 MB. Yesterday, the high-quality version was downloaded four times as often as the standard version. The total size downloaded for the two versions was 4536 MB. How many downloads of the standard version were there?" "The size of the standard version is 2.5 megabytes (MB)" = 2.5 "The size of the high-quality version is 4.1 MB." = 4.1 "Yesterday, the high-quality version was downloaded four times as often as the standard version." = 4.1(4x) + 2.5(x) = _ "The total size downloaded for the two versions was 4,536 MB. How many downloads of the standard version were there?" = 2.5(x) + 4.1(4x) = 4,536 That would be my conclusion. I hope this helps in someway! Have a great day! ;) {---Ghostgate---}

jimthompson5910 (jim_thompson5910):

after that you can replace the y with 4x (since y = 4x) 2.5x+4.1y = 4536 2.5x+4.1(4x) = 4536 I'll let you solve

OpenStudy (unofficialllyy):

I got 240

jimthompson5910 (jim_thompson5910):

x = 240 is correct

OpenStudy (ghostgate):

In the end jim and I came up with the same equation as the answer. Lol - Nice work jim. ;)

OpenStudy (unofficialllyy):

Thank you guys!

jimthompson5910 (jim_thompson5910):

then use y = 4x and x = 240 to find y